OpenStudy (anonymous):

z1= -2-4i z2= 5-2i What is the product of z1 and its conjugate?

OpenStudy (anonymous):

The i means imaginary number.

OpenStudy (anonymous):

conjugate of z1 is (-2+4i) z1* z1conjugate = (-2-4i)(-2+4i) = 4 + 16=20